OpenStudy (anonymous):

Solve. x2 + 10x = –2

OpenStudy (rulnick):

Is that x^2 + 10x = -2 ?

OpenStudy (anonymous):

yes!

OpenStudy (rulnick):

x^2 + 10x + 2 = 0 x = (-10 +/- sqrt(100-8)) / 2 = -5 +/- sqrt(23)

OpenStudy (rulnick):

approx. -9.8 and -0.2

OpenStudy (anonymous):

First arrange it so that it is equal to 0: \[x^{2} + 10x + 2 = 0\] Then use the quadratic equation, as rulnick did:\[x = \frac{-b \pm \sqrt{b^{2} - 4ac}}{2a}\] Where a is the coefficient of the first term (1), b is the coefficient of the 2nd term (10), and c and the coefficient of the 3rd term (2)

OpenStudy (anonymous):

Thank you guys!

OpenStudy (agreene):

Or, if you don't know/ don't like the quadratic equation: x^2+10x=-2 add 25 to both sides x^2+10x+25=23 factor the left (x+5)^2 =23 take the squareroot of both sides |x+5|=sqrt(23) you then eliminate the absolute value, giving the answers: \[x=-5\pm\sqrt{23}\] this method is known as completing the square.